# Inferring Thunderstorm Occurrence from Vertical Profiles of Convection-Permitting Simulations: Physical Insights from a Physical Deep Learning Model

### v2

@@ Abstract
Thunderstorms have significant social and economic impacts due to heavy precipitation, hail, lightning, and strong winds, necessitating reliable forecasts. Thunderstorm forecasts based on numerical weather prediction (NWP) often rely on single-level surrogate predictors, like convective available potential energy and convective inhibition, derived from vertical profiles of three-dimensional atmospheric variables. In this study, we develop SALAMA 1D, a deep neural network which directly infers the probability of thunderstorm occurrence from vertical profiles of ten atmospheric variables, bypassing single-level predictors. By training the model on convection-permitting NWP forecasts, we allow SALAMA 1D to flexibly identify convective patterns, with the goal of enhancing forecast accuracy. The model’s architecture is physically motivated: sparse connections encourage interactions at similar height levels while keeping model size and inference times computationally efficient, whereas a shuffling mechanism prevents the model from learning non-physical patterns tied to the vertical grid. SALAMA 1D is trained over Central Europe with lightning observations as the ground truth. Comparative analysis against a baseline machine learning model that uses single-level predictors shows SALAMA 1D’s superior skill across various metrics and lead times of up to at least 11 hours. Moreover, expanding the archive of forecasts from which training examples are sampled improves skill, even when training set size remains constant. Finally, a sensitivity analysis using saliency maps indicates that our model relies on physically interpretable patterns consistent with established theoretical understanding, primarily focusing on convection initiation and refining its skill by checking for ongoing convection in the NWP model.
@@ 1 Introduction
The predictors that have been used by ML models for inferring thunderstorm occurrence are single-level variables; i.e., they assign a single value to each horizontal grid point on the NWP model domain. Most single-level predictors can be derived from the vertical profiles of three-dimensional meteorological variables. For example, CAPE is determined by the vertical profiles of pressure, temperature, and specific humidity (Markowski and Richardson, 2010). There are notable exceptions, e.g., orography, or solar radiation, which are possible sources of lift which forecasters consider when predicting convection initiation. In what follows, however, we restrict ourselves to single-level predictors that derive from vertical profiles. Importantly, the information contained in a set of such single-level predictors is inherently present in the corresponding vertical profiles, though in a more complex, encoded form. Consequently, an ML model trained directly on vertical profiles should, at a minimum, match the performance of one trained on single-level predictors. Given that a model based on vertical profiles has greater flexibility to detect patterns in NWP data, we hypothesize that such a model would outperform one using only single-level predictors—this is the primary focus of the present study.
@@
A major issue of ML models concerns the interpretability of their output (Dramsch et al., 2025). In this work, we conduct a sensitivity analysis using saliency maps to explore how the model processes the input vertical profiles. As this analysis will reveal, our model rediscovers patterns that align with established theoretical understandings of thunderstorm development and occurrence.
@@ 2 Data
Our objective is to train an ML model—SALAMA 1D—for inferring the conditional probability of thunderstorm occurrence given an array of meteorological variables at height levels. As a baseline for comparison, we will also train a separate ML model (SALAMA 0D, Section 33.3) which infers the probability of thunderstorm occurrence from single-level meteorological variables. The ML task is known as binary classification. ML models for this task are trained with data sets of examples. In our case, an example is a tuple , where the label denotes the class to which the example belongs (class 1: thunderstorm occurrence, class 0: no thunderstorm occurrence). We first collect an archive of NWP forecasts of (Section 22.1) and the corresponding labels from lightning observations (Section 22.2). We then randomly draw examples from the archive to compile data sets for training, validation, and testing (Section 22.3).
@@ 2.1 NWP data
The operational runs of ICON-D2-EPS are initialized daily every , starting at 0000 UTC, and produce forecasts with a time resolution of . For each full hour of the day, we collect only the latest (and, therefore, most accurate) available forecast, which is at most old. We gather forecasts for several summer months in 2021 (June, July, August), 2022 (May, June, July), and 2023 (July, August). ICON-D2-EPS comprises 20 ensemble members, which reflect the NWP uncertainty in the initial conditions, model error, and boundary conditions (Reinert et al., 2020). We collect the forecasts for all members of the ensemble system.
@@ 2.2 Lightning observations
where denotes the great-circle distance between and on a perfect sphere of the Earth with a radius of , as assumed in the ICON-D2-EPS model (Reinert et al., 2020). The spatial and temporal thresholds used in this study read and . NWP models struggle with accurately predicting the correct location and timing of convection (Roberts and Lean, 2008); therefore, grid points with a positive class label do not reliably coincide with convective patterns in the NWP model. We revisit this issue in Section 44.2.
@@ 4.1 Model comparison
Raw NWP output structures have a smaller horizontal extension than the lightning regions in both cases, resulting in a considerable number of missed pixels even when the NWP model produces convection within a lightning region. The ML models tend to correct the size of the convective structures, with SALAMA 1D-2022 producing the least false alarms. Remarkably, the SALAMA 1D models can also produce high-probability output when lightning occurred but no convection has been triggered in the NWP model, as can be seen for the lightning regions over France for case A, which suggests that our ML models, SALAMA 1D-2022 in particular, may be able to correct for NWP model biases.
@@ 4.2 Interpretability study
The first column of panels in Fig. 7 shows temperature and dew-point temperature for the two percentiles. The top percentile troposphere displays more moisture than the bottom percentile, in particular in 2- height. This is consistent with the documented importance of moisture for thunderstorm development, as the buoyancy of rising air parcels is otherwise reduced by dry-air entrainment (Zhang et al., 2003; Peters et al., 2023; Marquis et al., 2023). In the second column in Fig. 7, we show average profiles of mixing ratios of cloud water (QC), cloud ice (QI), and graupel (QG). This suggests that the model uses non-vanishing profiles of QI and QG to discriminate between the top and the bottom percentile. For high-probability samples, ice particle content peaks close to the tropopause at a height of 10-, consistent with measurements of vertical hydrometeor distributions (e.g., Vivekanandan et al., 1999; Hubbert et al., 2018). The third column in Fig. 7 displays the average profiles of cloud cover (CLC). In general, our model associates non-vanishing CLC with a high probability of thunderstorm occurrence. In particular, close to the tropopause, CLC tends to be . This is consistent with anvil cloud top levels (Markowski and Richardson, 2010). In the fourth column in Fig. 7, we show vertical profiles of the three wind components U, V, and W. It is noteworthy that high-probability samples tend to have southwesterly wind profiles, whereas samples from the bottom percentile display northwesterly winds. This is consistent with studies on the typical propagation direction of thunderstorms in Central Europe (Hagen and Finke, 1999). On the other hand, vertical profiles of W vanish for both percentiles. We presume that due to convection displacement errors in the NWP model, the updraft regions within simulated deep convection rarely matches observed lightning observations. Therefore, SALAMA 1D may have learned not to rely on W for inferring thunderstorm occurrence. The last column of Fig. 7 shows the average profiles of pressure. Both profiles appear to be essentially hydrostatic. Surface pressure tends to be lower for the top percentile than for the bottom percentile.
@@
While saliency varies from one sample to another and can be used to interpret individual predictions, we propose averaging over the top probability percentile to obtain more robust insight on the input fields and height levels which contribute most to high-probability output. To this end, we compute , where the angle brackets denote averaging over the top percentile samples. We take the absolute value (absolute saliency), as we consider feature importance to be linked to the (sample-averaged) intensity of the ML model’s linear response, irrespective of whether this response is positive or negative. In what follows, we refer to simply as saliency, i.e., with sample-averaging and taking the absolute value being implicitly implied unless stated otherwise. The resulting saliency maps for the two percentiles are shown in Fig. 8. Note that average saliencies of the different fields are stacked on top of each other. As a consequence, the saliency envelope quantifies how much individual height levels affect the model outcome.
@@
Next, we turn to the mid-level peak. Apart from horizontal wind velocity, the mid-level peak receives most contributions from temperature. In addition, pressure saliency is non-vanishing only at the mid-level peak and near the surface. To understand this, note that our model is not informed about the height of individual vertical levels; rather, these levels are randomly shuffled (Section 33.1). Thus, our model can reliably infer heights levels—and in particular level spacings—only from pressure, which monotonously decreases with height (Fig. 7). Therefore, we expect pressure saliency to be a proxy for how much the model relies on vertical gradients. Since pressure saliency contributes to the mid-level peak and temperature saliency is high, we hypothesize that our model reconstructs mid-level temperature lapse rates. This is supported by the fact that the mid-level peak is bounded by the LFC and the EL of parcels lifted from the mixed layer, meaning that such parcels are buoyant for height levels in the vicinity of the mid-level peak. Positive buoyancy occurring in a conditionally unstable troposphere is known to be a crucial ingredient for thunderstorm development (Doswell et al., 1996) and constitutes the basis for several traditional thunderstorm predictors, such as convective available potential energy (CAPE). To test whether SALAMA 1D considers mid-level lapse rates, we show in Fig. 9a the distribution of 500- lapse rates for the top and bottom probability percentile. Indeed, essentially all high-probability samples are associated with conditionally unstable mid-level lapse rates, whereas the bottom percentile distribution extends further into absolutely stable mid-level lapse rates. Correspondingly, the top percentile profiles exhibit significantly more CAPE than those in the bottom percentile. (Fig. 9c). On the other hand, the distributions of the two percentiles show a significant overlap, which implies that considering mid-level lapse rates alone is not sufficient to infer a high probability of thunderstorm occurrence. As pressure saliency is non-vanishing also near the surface, we show in Fig. 9b distributions of near-surface (10-) lapse rates. Indeed, the distributions differ for the two percentiles. In contrast to before, near-surface lapse rates for the top percentile tend to be lower, and mostly absolutely stable, whereas lapse rates for the bottom percentile are mostly conditionally unstable. This is consistent with finding higher convective inhibition (CIN) in the top percentile than in the bottom percentile (Fig. 9d). This illustrates how the presence of CIN allows for CAPE to build up, which in turn makes convection initiation more likely (Carlson et al., 1983; Tuckman et al., 2023).
@@
horizontal wind direction,
mid-level and near-surface temperature lapse rates, or
low-level moisture.
We refer to these patterns as mesoscale since the underlying fields, such as temperature and pressure, vary more slowly in the horizontal than the fine-grained hydrometeor variables do. With the exception of horizontal wind direction, these patterns tend to be more characteristic of convective environments, in which convection is likely to be initiated, than of ongoing convection. In consistence with this, the top percentile samples show higher values of CAPE and CIN than the bottom percentiles do. Conversely, we identify a category of sub-mesoscale patterns, which relate to
@@
The sub-mesoscale patterns indicate ongoing convection. It is noteworthy that in terms of saliency, our model seems to rely less on sub-mesoscale patterns than on mesoscale patterns. As predicting the correct location and timing of convection remains challenging even for convection-permitting models (Roberts and Lean, 2008), SALAMA 1D may have learned to be cautious about sub-mesoscale patterns if they do not reliably match lightning observations. As a result, our model—on average—places greater emphasis (as indicated by saliency) on mesoscale patterns, while sub-mesoscale patterns serve to refine its skill. In practice, this also implies that SALAMA 1D is not strictly reliant on convection being produced in the NWP model. Instead, even if only some of the ingredients for convection are present, our model assigns a likely moderate but still well-calibrated probability of thunderstorm occurrence. In consequence, we expect some skill at correcting NWP output, as already suggested by the case studies in Fig. 4.
@@ 5 Discussion and conclusion
A sensitivity analysis based on saliency maps revealed that many learned patterns are physically interpretable and indicated that our model significantly relies on mesoscale patterns: For instance, our results suggested that SALAMA 1D has learned the climatological propagation direction of thunderstorms in the study region and involves temperature, pressure and moisture profiles to infer atmospheric instability and, hence, the potential for convective initiation. To a lesser extent, the model relies on more fine-grained structures characteristic of ongoing convection, such as ice particle content near the tropopause and cloud cover. We hypothesized that the preference for mesoscale patterns arises because of their higher reliability for identifying thunderstorm occurrence, as fine-grained patterns may not consistently align with lightning observations due to the convection-permitting NWP model struggling to accurately predict the location and timing of convection. To encourage the use of fine-grained structures, it may be beneficial to train the model in such a way that it can process all ensemble members simultaneously in one forward pass. By analyzing the number of members displaying fine-grained signs of thunderstorm activity, the model could learn the degree to which it can trust the convection produced by NWP. Furthermore, one could adapt the model in such a way that horizontally extended input, or several forecast times around the target time, are processed. This would allow for improvement on correcting location and timing errors of convection.

### v3

@@ Abstract
Thunderstorms have significant social and economic impacts due to heavy precipitation, hail, lightning, and strong winds, necessitating reliable forecasts. Thunderstorm forecasts based on numerical weather prediction (NWP) often rely on single-level surrogate predictors, like convective available potential energy and convective inhibition, derived from vertical profiles of three-dimensional atmospheric variables. In this study, we develop SALAMA 1D, a deep neural network which directly infers the probability of thunderstorm occurrence from vertical profiles of ten atmospheric variables, bypassing single-level predictors. By training the model on convection-permitting NWP forecasts, we allow SALAMA 1D to flexibly identify convective patterns, with the goal of enhancing forecast accuracy. The model’s architecture is physically motivated: sparse connections encourage interactions at similar height levels while keeping model size and inference times computationally efficient, whereas a shuffling mechanism prevents the model from learning non-physical patterns tied to the vertical grid. SALAMA 1D is trained over Central Europe with lightning observations as the ground truth. Comparative analysis against a baseline machine learning model that uses single-level predictors shows SALAMA 1D’s superior skill across various metrics and lead times of up to at least 11 hours. Moreover, expanding the archive of forecasts from which training examples are sampled improves skill, even when training set size remains constant. Finally, a sensitivity analysis using saliency maps indicates that our model relies on physically interpretable patterns consistent with established theoretical understanding, such as ice particle content near the tropopause, cloud cover, conditional instability, and low-level moisture.
@@ 1 Introduction
The predictors that have been used by ML models for inferring thunderstorm occurrence are single-level variables; i.e., they assign a single value to each horizontal grid point on the NWP model domain. Most single-level predictors can be derived from the vertical profiles of three-dimensional meteorological variables. For example, CAPE is determined by the vertical profiles of pressure, temperature, and specific humidity (Markowski and Richardson, 2010). There are additional processes and effects outside the scope of vertical profiles, e.g., orography, or solar radiation, which are possible sources of lift which forecasters consider when predicting convection initiation. In what follows, however, we restrict ourselves to single-level predictors that are derived from vertical profiles. Importantly, the information contained in a set of such single-level predictors is inherently present in the corresponding vertical profiles, though in a more complex, encoded form. Consequently, an ML model trained directly on vertical profiles should, at a minimum, match the performance of one trained on single-level predictors. Given that a model based on vertical profiles has greater flexibility to detect patterns in NWP data, we hypothesize that such a model would outperform one using only single-level predictors—this is the primary focus of the present study.
@@
A major issue of ML models concerns the interpretability of their output (Flora et al., 2024; Yang et al., 2024; Dramsch et al., 2025). In this work, we conduct a sensitivity analysis using saliency maps to explore how the model processes the input vertical profiles. As this analysis will reveal, our model rediscovers patterns that align with established theoretical understandings of thunderstorm development and occurrence.
@@ 2 Data
Our objective is to train an ML model—SALAMA 1D—for inferring the conditional probability of thunderstorm occurrence given an array of meteorological variables at height levels. As a baseline for comparison, we will also train a separate ML model (SALAMA 0D, Section 33.3) which infers the probability of thunderstorm occurrence from single-level meteorological variables. The ML task is known as binary classification. ML models for this task are trained with data sets of examples. In our case, an example is a tuple , where the label denotes the class to which the sample belongs (class 1: thunderstorm occurrence, class 0: no thunderstorm occurrence). We first collect an archive of NWP forecasts of (Section 22.1) and the corresponding labels from lightning observations (Section 22.2). We then randomly draw examples from the archive to compile data sets for training, validation, and testing (Section 22.3).
@@ 2.1 NWP data
The operational runs of ICON-D2-EPS are initialized eight times daily, starting at 0000 UTC, and produce forecasts with a time resolution of . For each full hour of the day, we collect only the latest (and, therefore, most accurate) available forecast, which is at most old. We gather forecasts for several summer months in 2021 (June, July, August), 2022 (May, June, July), and 2023 (July, August). ICON-D2-EPS comprises 20 ensemble members, which reflect the NWP uncertainty in the initial conditions, model error, and boundary conditions (Reinert et al., 2020). We collect the forecasts for all members of the ensemble system.
@@ 2.2 Lightning observations
where denotes the great-circle distance between and on a perfect sphere of the Earth with a radius of , as assumed in the ICON-D2-EPS model (Reinert et al., 2020). The spatial and temporal thresholds used in this study read and .
@@ 4.1 Model comparison
The ML models, overall, align with raw NWP structures, with highest ML probability output being collocated with a high likelihood of exceeding . On the other hand, the ML models tend to correct the areal size of simulated convection, with SALAMA 1D-2022 producing the least false alarms. Remarkably, the SALAMA 1D models can also produce high-probability output when lightning occurred but no convection has been triggered in the NWP model, as can be seen for the lightning regions over France for case A, which suggests that our ML models, SALAMA 1D-2022 in particular, may be able to correct for NWP model biases.
@@ 4.2 Interpretability study
The first column of panels in Fig. 7 shows temperature and dew-point temperature for the two percentiles. The top percentile troposphere displays more moisture than the bottom percentile, in particular in 2- height. This is consistent with the documented importance of moisture for thunderstorm development, as the buoyancy of rising air parcels is otherwise reduced by dry-air entrainment (Zhang et al., 2003; Peters et al., 2023; Marquis et al., 2023). In the second column in Fig. 7, we show average profiles of mixing ratios of cloud water (QC), cloud ice (QI), and graupel (QG). This suggests that the model uses non-vanishing profiles of QI and QG to discriminate between the top and the bottom percentile. For high-probability samples, ice particle content peaks close to the tropopause at a height of 10-, consistent with measurements of vertical hydrometeor distributions (e.g., Vivekanandan et al., 1999; Hubbert et al., 2018). The third column in Fig. 7 displays the average profiles of cloud cover (CLC). In general, our model associates non-vanishing CLC with a high probability of thunderstorm occurrence. In particular, close to the tropopause, CLC tends to be . This is consistent with anvil cloud top levels (Markowski and Richardson, 2010). In the fourth column in Fig. 7, we show vertical profiles of the three wind components U, V, and W. It is noteworthy that high-probability samples tend to have southwesterly wind profiles, whereas samples from the bottom percentile display northwesterly winds. This is consistent with studies on the typical propagation direction of thunderstorms in Central Europe (Hagen and Finke, 1999). On the other hand, vertical profiles of W vanish for both percentiles. We presume that due to convection displacement errors in the NWP model, the updraft regions within simulated deep convection rarely match observed lightning observations. Therefore, SALAMA 1D may have learned not to rely on W for inferring thunderstorm occurrence. The last column of Fig. 7 shows the average profiles of pressure. Both profiles appear to be essentially hydrostatic. Surface pressure tends to be lower for the top percentile than for the bottom percentile.
@@
While saliency varies from one sample to another and can be used to interpret individual predictions, we propose averaging over the top probability percentile to obtain more robust insight on the input fields and height levels which contribute most to high-probability output. To this end, we compute , where the angle brackets denote averaging over the top percentile samples. We take the absolute value (absolute saliency), as we consider feature importance to be linked to the (sample-averaged) intensity of the ML model’s linear response, irrespectively of whether this response is positive or negative. In what follows, we refer to simply as saliency, i.e., with sample-averaging and taking the absolute value being implicitly implied unless stated otherwise. The resulting saliency maps for the two percentiles are shown in Fig. 8. Note that average saliencies of the different fields are stacked on top of each other. As a consequence, the saliency envelope quantifies how much individual height levels affect the model outcome.
@@
Next, we turn to the mid-level peak. Apart from horizontal wind velocity, the mid-level peak receives most contributions from temperature. This may be partly due to convection feeding back to the temperature field. However, since pressure saliency is non-vanishing only at the mid-level peak and near the surface, we hypothesize that the ML model also reconstructs lapse rates. To understand this, note that our model is not informed about the height of individual vertical levels; rather, these levels are randomly shuffled (Section 33.1). Thus, our model can reliably infer heights levels—and in particular level spacings—only from pressure, which monotonously decreases with height (Fig. 7). Therefore, we expect pressure saliency to be a proxy for how much the model relies on vertical gradients. Finally, as pressure saliency contributes to the mid-level peak and temperature saliency is high, we conjecture that our model reconstructs mid-level lapse rates. This is supported by the fact that the mid-level peak is bounded by the LFC and the EL of parcels lifted from the mixed layer, meaning that such parcels are buoyant for height levels in the vicinity of the mid-level peak. Positive buoyancy occurring in a conditionally unstable troposphere is known to be a crucial ingredient for thunderstorm development (Doswell et al., 1996) and constitutes the basis for several traditional thunderstorm predictors, such as convective available potential energy (CAPE).
To test whether SALAMA 1D considers mid-level lapse rates, we show in Fig. 9a the distribution of 500- lapse rates for the top and bottom probability percentile. Indeed, essentially all high-probability samples are associated with conditionally unstable mid-level lapse rates, whereas the bottom percentile distribution extends further into absolutely stable mid-level lapse rates. On the other hand, the distributions of the two percentiles show a significant overlap, which implies that considering mid-level lapse rates alone is not sufficient to infer a high probability of thunderstorm occurrence.
Complementarily to mid-level lapse rates, we show the distributions of CAPE for the two percentiles in Fig. 9c. Most samples in both percentiles have a low value of CAPE, with around of the high-probability samples and of the low-probability samples falling into the lowest bin. Nevertheless, the top percentile distribution has a longer tail, towards higher CAPE values, than the bottom percentile distribution does. The low CAPE values of the bottom percentile samples are consistent with conditional instability failing to develop. As for the top percentile samples, we expect CAPE to be considerably reduced inside the core of a convective cell, whereas some CAPE is expected to remain in lightning regions of less intense precipitation, producing the long tail of the top percentile distribution.
As pressure saliency is non-vanishing also near the surface, we show in Fig. 9b distributions of near-surface (10-) lapse rates. Indeed, the distributions differ for the two percentiles. In contrast to the mid-level peak, near-surface lapse rates for the top percentile tend to be lower, and mostly absolutely stable, whereas lapse rates for the bottom percentile are mostly conditionally unstable. Figure 9d shows the corresponding distributions of convective inhibition (CIN). While most samples of both percentiles fall into the lowest bin, the top percentile distribution has a longer tail towards higher values of CIN. This might seem surprising as this result seems to suggest that more stable low-level lapse rates (or: higher values of CIN) favor thunderstorm occurrence. However, the reduction of CIN is only conducive to thunderstorm development if sufficient CAPE has been able to build up beforehand. The bottom-percentile samples are essentially from environments with low CAPE and low CIN, which suggests that in these cases, the constantly low CIN (due to, e.g., continuous mixing of low-level air) caused any instabilities aloft to be released prematurely, preventing CAPE from building up, or thunderstorms from forming (Carlson et al., 1983; Tuckman et al., 2023). As for the top-percentile samples, CIN is expected to be mostly removed within storm centers, while some CIN can be expected to remain in the less intense regions of convective precipitation or for samples with simulated convection occurring nearby or in the near future. This effect likely produces the longer tail of the top-percentile distribution of CIN.
@@
These patterns pertain to regions within ongoing convection, in which precipitation is most intense. In contrast, we identify a category of patterns related to
horizontal wind direction,
mid-level and near-surface lapse rates, or
low-level moisture.
We refer to the latter patterns as mesoscale since the underlying fields, such as temperature and pressure, vary more slowly in the horizontal than the fine-grained hydrometeor variables do. Conversely, we refer to the former patterns as sub-mesoscale. Mesoscale patterns in the above sense tend to be characteristic of regions with sufficient distance to the convective cores such that precipitation is less intense and some CAPE remains to be released. SALAMA 1D is sensitive to both categories in order to identify thunderstorm occurrence. While sub-mesoscale patterns are useful for the identification of convective storm centers, the sensitivity to mesoscale patterns could explain the ML model’s skill at increasing the areas of simulated convection observed in the case study (Fig. 4), namely by accounting for modestly-precipitating regions with low-level moisture and left-over CAPE.
@@ 5 Discussion and conclusion
A sensitivity analysis based on saliency maps revealed that many learned patterns are physically interpretable. For instance, our results suggested that SALAMA 1D has learned the climatological propagation direction of thunderstorms in the study region and relies on fine-grained (sub-mesoscale) structures, such as ice particle content near the tropopause, or cloud cover, to identify high-precipitation regions of ongoing convection. Conversely, mesoscale patterns related to atmospheric instability and moisture are used, possibly to account for regions with less intense precipitation and left-over CAPE. We hypothesized that mesoscale patterns are instrumental in correcting the areal size of simulated convection.
To improve the ML model’s capability of correcting for NWP-related biases, it may be beneficial to adapt the model in such a way that horizontally extended input, or several forecast times around the target time, are processed. This would allow for improvement on correcting location and timing errors of convection. Furthermore, one could train the model in such a way that it can process all ensemble members simultaneously in one forward pass, which would enable the ML model to account for the NWP forecast uncertainty.
@@ Acknowledgements.
- Flora et al. (2024) Flora, M. L., C. K. Potvin, A. McGovern, and S. Handler, 2024: A machine learning explainability tutorial for atmospheric sciences. Artificial Intelligence for the Earth Systems, 3 (1), e230 018, 10.1175/AIES-D-23-0018.1.
@@
- Yang et al. (2024) Yang, R., and Coauthors, 2024: Interpretable machine learning for weather and climate prediction: A review. Atmospheric Environment, 338, 120 797, https://doi.org/10.1016/j.atmosenv.2024.120797.